The annual increase in height of cedar trees is believed to be distributed uniformly between three and fourteen inches. Find the probability that a randomly selected cedar tree will grow more than 6 inches in a given year. Round your answer to four decimal places, if necessary.
Answer #1

Solution :

Given that,

a = 3

b = 14

P(x > c) = (b - c) / (b - a)

P(x > 6) = (14 - 6) / (14 - 3)

= 8 / 11 = 0.7273

Probability = 0.7273